HCl molecule shows an absorption at 2886 cm^-1. Determine its force constant and maximum displacements (change in inter nuclear distance) for v=0,1,2,3
1
Expert's answer
2019-05-27T06:09:24-0400

Solution.

"\mu = \frac{m(H) \times m(Cl)}{m(H) + m(Cl)}"

"\mu = \frac{1 \times 35.5}{1 + 35.5}"

"\mu = 0.97 a.m.u."

"\mu = 0.97 \times 1.66*10^{-27}"

"\mu = 1.61*10^{-27} kg"

"K = \mu \times (\frac{2 \times \pi \times c}{\lambda})^2"

"\lambda = \frac{1}{2886} * 10^{-7} = 3465 nm"

"K = 1.61*10^{-27} * (\frac{2*3.14*3*10^8}{3465*10^{-9}})^2"

"K = 475.97"

"E=\frac{h*c}{\lambda}"

"J = \frac{h^2}{\Delta E}"

"J = 1.21*10^{-48}"

"d = \sqrt{(\frac{J}{\mu})}"

"d = 2.74*10^{-11}"

Answer:

K = 475.97 N/m

d = 2.74*10^(-11) m